OpenStudy (anonymous):

Integrate x^2+2x+3/x^3+3x^2+9x

OpenStudy (anonymous):

try a u - sub

OpenStudy (anonymous):

\[y=x^3+3x^2+9x\\ du=(3x^2+6x+9)dx\\ \frac{du}{3}=(x^2+2x+3)dx\] pretty well cooked this one

OpenStudy (misty1212):

HI!!

OpenStudy (misty1212):

you got this after the u substitution? you are going to end up with \[\frac{1}{3}\int \frac{du}{u}\] which you solve in one step `

OpenStudy (solomonzelman):

sate, just that you wrote y=... and then du=... very good approach though

OpenStudy (anonymous):

Thank you!

OpenStudy (solomonzelman):

I mean, \(\large\color{slate}{ u=x^3+3x^2+9x }\) \(\large\color{slate}{ du=(3x^2+6x+9)dx }\) \(\large\color{slate}{ (1/3)du=(x^2+x+3)dx }\)

OpenStudy (solomonzelman):

(just a little variable err)

OpenStudy (solomonzelman):

camila, and then make sure to switch back to x, as you are done

OpenStudy (anonymous):

If you're feeling particularly ambitious you can also split up the given expression into partial fractions: \[\frac{x^2+2x+3}{x^3+3x^2+9x}=\frac{x^2+2x+3}{x(x^2+3x+9)}=\frac{A}{x}+\frac{Bx+C}{x^2+3x+9}\]